OpenStudy (anonymous):

Please help, 2x + 4y +3z=6 3x + 5y +6z=3 2x +3y +4z+8 I would really appreciate any help

OpenStudy (anonymous):

given equations are 2x+4y+3z = 6 3x+5y+6z = 3 2x+3y+4z = 8 now step one multiply eq 1 with 3 we will get 3(2x+4y+3z = 6) 6x+12y+9z = 18